# The Color-Clinical Decoupling: Why Perceptual Calibration Fails Clinical Biomarkers in Smartphone Dermatology

### v1

@@ Abstract
Methods: We analyzed 43,425 images from 965 Korean subjects (Fitzpatrick III–IV) across DSLR, tablet, and smartphone devices using the AI Hub Korean Skin Condition Measurement dataset. Linear Color Correction Matrix (CCM) normalization—the industry standard—was applied to align consumer devices to a DSLR reference. Performance was assessed using CIEDE2000 color differences ($\Delta E$) and Intraclass Correlation Coefficients (ICC) for Melanin Index, Erythema Index, and Individual Typology Angle (ITA).
@@ 1 Introduction
The standard solution—linear Color Correction Matrix (CCM) normalization—transforms device RGB values toward a reference standard by minimizing the Euclidean distance in color space.[Hong2001, Cheung2004] This approach optimizes for perceptual color fidelity, as measured by the CIEDE2000 color difference formula ($\Delta E_{00}$).[Sharma2005, Luo2001] The clinical acceptability threshold of $\Delta E<2$ corresponds to the just-noticeable difference in human color perception.[Melgosa2001]
@@
Prior research has validated smartphone telemedicine for specific conditions,[Lamel2012, VanoGalvan2015] developed multi-device domain adaptation for other imaging modalities,[Chen2019, Tellez2019, Guan2021] and established RGB-based chromophore quantification methods.[Tsumura1999, Stamatas2004] However, these studies predominantly utilized datasets composed of Fitzpatrick I–II skin types,[Daneshjou2022, Groh2021] and none systematically tested whether colorimetric calibration guarantees clinical biomarker preservation.
This study addresses this critical gap. Using 43,425 images from 965 Korean subjects (Fitzpatrick III–IV), we test the hypothesis that standard colorimetric calibration ensures clinical reliability. Our investigation reveals three key findings:
@@
    Color-Clinical Decoupling: We introduce and systematically investigate “color-clinical decoupling”—the phenomenon where perceptually accurate calibration ($\Delta E<2.3$) fails to preserve clinical biomarkers. While prior work has identified ITA measurement limitations including sensitivity to illuminants,[Jung2024] this is the first systematic demonstration that standard colorimetric calibration creates false clinical confidence, with ITA showing poor agreement (ICC =0.40=0.40) despite excellent color fidelity. We provide mathematical derivation explaining this phenomenon through the differential sensitivity of ratio-based formulas to channel-specific errors.
@@ 2.1 Optical Properties of Human Skin
Human skin functions as a multilayered turbid medium whose appearance is governed by the interplay of absorption and scattering within the epidermis and dermis.[Tuchin2015, Jacques2013] The two principal chromophores are melanin, concentrated in the basal layer of the epidermis, and hemoglobin, distributed within the dermal vasculature.[Anderson1981] Melanin exhibits broadband absorption across the visible spectrum, decreasing monotonically from 300 to 700 nm. Oxyhemoglobin displays characteristic absorption peaks at 542 nm and 577 nm, and blood stasis patterns further modulate perceived skin pigmentation.[Zonios2001, Stamatas2008]
The spectral reflectance of skin results from a complex light transport process involving specular reflection at the stratum corneum, volumetric scattering within the dermis, and wavelength-dependent absorption at varying depths.[Meglinski2002, Baranoski2004] RGB sensors integrate this continuous spectrum into three broad channels, irreversibly discarding the fine spectral detail required to deconvolve individual chromophore contributions—a phenomenon known as metamerism.[Nishidate2004]
@@ 2.2 Device-Dependent Color Rendering
Modern digital cameras exhibit substantial variability in color reproduction due to differences in sensor design and ISP pipelines.[Karaimer2016] Recent assessments of smartphone cameras for digital health applications have confirmed persistent color accuracy limitations across consumer devices.[Yoon2020] Three primary sources of device dependence exist: (1) spectral sensitivity functions of color filter arrays vary across manufacturers;[Hong2001, Chakrabarti2014] (2) ISP pipelines apply nonlinear transformations including gamma correction, tone mapping, and “memory color” optimization;[Karaimer2016, Cheung2004] and (3) the sRGB color space provides limited gamut and introduces quantization errors in dark tones.[Finlayson2015]
@@ 2.3 Clinical Skin Color Indices
Quantitative dermatology employs several derived indices to characterize skin pigmentation. The Individual Typology Angle (ITA) provides an objective correlate of Fitzpatrick skin phototype.[Chardon1991, Wilkes2015, DelBino2006] The Melanin Index, computed from $L^{*}$ using logarithmic transformations, correlates with epidermal melanin concentration.[Chardon1991, Stamatas2004] The Erythema Index, derived from the $a^{*}$ channel, reflects dermal hemoglobin content.[Fullerton1996, Takiwaki1998] Critically, while device validation studies have assessed inter-device agreement for individual indices,[Yoon2020] no prior work has systematically compared how different clinical metrics—with their distinct mathematical structures—respond to identical calibration procedures. This gap is significant: a calibration optimized for one index may fail another.
@@ 2.4 The Implicit Assumption of Colorimetric Calibration
For typical skin colors in our dataset ($L^{*}\approx 75$, $b^{*}\approx 15$), equation (5) yields $|\partial\text{ITA}/\partial b^{*}|\approx 0.04$ rad/unit $\approx 2.3$°/unit. A seemingly small 1-unit error in $b^{*}$ produces a 2–3° shift in ITA—sufficient to cross Fitzpatrick skin type boundaries—while contributing only $\sim$1 to the $\Delta E$ calculation. Notably, perceptibility thresholds differ across CIELAB axes: thresholds of approximately 1.0 $\Delta E$ units apply to $L^{*}$ and $a^{*}$, but $\sim$2.6 units for $b^{*}$,[LindseyWee2007] meaning $b^{*}$ errors are both less perceptible and more clinically consequential.
The $b^{*}$ channel is derived primarily from the blue sensor, which exhibits the lowest quantum efficiency among RGB channels (approximately 30% of green channel efficiency).[Ly2020] Shot noise scales as $\sigma\propto 1/\sqrt{QE}$, making blue the noisiest channel by a factor of $\sim$1.8. This noise propagates through sRGB-to-CIELAB transformation and is amplified by the ITA formula’s sensitivity to $b^{*}$.
@@ 2.5 Dataset Bias in Dermatological AI
While underrepresentation of darker skin types in dermatological AI is well-documented,[Kinyanjui2020, Groh2021, Adamson2018] our dataset of 43,425 images from 965 Korean subjects (Fitzpatrick III–IV) represents the largest cohort specifically designed to evaluate color-clinical decoupling in this underrepresented population, providing systematic characterization of device-dependent color rendering.
@@ 3.1.2 Imaging Devices
    DSLR (Reference): Professional-grade camera under controlled dark room lighting (33,774 samples, 7 angles)
@@
    Tablet: Consumer-grade tablet camera (4,825 samples, 3 angles)
@@
    Smartphone: Consumer-grade smartphone camera (4,825 samples, 3 angles)
@@ 4.1 Raw Device Performance: Clinical Unusability of Uncalibrated Images
Table 1 summarizes device characteristics. Consumer devices exhibited systematic underexposure relative to DSLR ($L^{*}=72.7$$77.6$ vs. $81.3$) and compressed $b^{*}$ values (8.968.9610.1210.12 vs. $17.59$).
@@
Figure 1: RGB and CIELAB distributions across devices showing systematic shifts in color rendering.
Figure 2: Bland-Altman plots for $L^{*}$, $a^{*}$, $b^{*}$. Largest bias in $b^{*}$ channel (DSLR vs smartphone: 8.30).
@@ 4.2 CCM Normalization: Colorimetric Success
Linear CCM substantially reduced color differences (Table 3): 77.4% improvement for tablet ($\Delta E$: 8.33 $\rightarrow$ 1.88) and 67.4% for smartphone (7.02 $\rightarrow$ 2.29). Post-normalization median values approached clinical acceptability (1.66 and 2.09 respectively).
@@
Figure 3: $\Delta E$ distributions before and after CCM calibration.
@@ 4.3 The Decoupling: Clinical Index Failure Despite Colorimetric Success
Despite effective colorimetric normalization, clinical index agreement varied dramatically (Table 4). While Melanin Index achieved good agreement (ICC =0.767=0.767), ITA showed poor agreement (ICC =0.402=0.402). This represents the central finding: perceptual color accuracy does not guarantee clinical biomarker reliability.
@@
Figure 4: Clinical index distributions across devices after CCM normalization.
Figure 5: Correlation plots for clinical indices. Note the scatter in ITA despite good Melanin Index correlation.
@@ 4.4 Sources of Variance: Implications for Calibration Protocol Design
ANOVA revealed a striking hierarchy of variance sources (Table 5). Facial anatomical region accounted for 25.2% of variance ($\eta^{2}=0.252$)—3.6 times greater than device hardware effects ($\eta^{2}=0.070$).
@@
Figure 6: $\Delta E$ distributions by facial region revealing substantial anatomical variation.
Figure 7: Stratified $\Delta E$ analysis: (A) Age, (B) Gender, (C) Skin type, (D) Facial region.
@@ 5.1 The Color-Clinical Decoupling Phenomenon
The central finding of this study is the non-linear decoupling between colorimetric accuracy and clinical biomarker reliability. Despite achieving near-clinical color fidelity ($\Delta E<2.3$), the ITA index exhibited poor inter-device agreement (ICC =0.40=0.40). This finding challenges the fundamental assumption underlying smartphone dermatology: that “looking correct” guarantees “diagnosing correctly.”
@@
The blue sensor’s low quantum efficiency ($\sim$30% of green) produces elevated shot noise that propagates directly to $b^{*}$. Linear CCM, optimized for overall color fidelity, cannot selectively suppress this critical error source. The result is “color-clinical decoupling”: colorimetric success masking clinical failure.
@@ 5.2 Why Single-Patch Calibration Fails
This “anatomical dominance” arises from multiple physiological sources. The T-zone (forehead and glabella) exhibits higher sebaceous gland density than cheek regions, resulting in elevated sebum secretion that alters surface refractive index.[Man2009, Ohshima2013] Sebum forms a thin lipid film that increases specular reflection and modifies the balance between surface gloss and diffuse reflectance.[Baranoski2004] Under directional illumination, this regional variation produces angle-dependent color shifts.
Facial skin thickness varies substantially: the periorbital region is among the thinnest (0.5–1.0 mm), while chin and forehead are thicker (1.5–2.0 mm).[Nouveau2016, Pioro2021] Since skin color arises from chromophore absorption at specific depths, thickness variations alter the optical path length and relative contributions of melanin versus hemoglobin.

### v2

@@ Abstract
Methods: We analyzed a total dataset of 43,424 images from 965 Korean subjects (Fitzpatrick III–IV) using the AI Hub Korean Skin Condition Measurement dataset. Cross-device comparisons utilized 14,475 matched front-facing images (4,825 per device) acquired across DSLR, tablet, and smartphone devices. Linear Color Correction Matrix (CCM) normalization—the industry standard—was applied to align consumer devices to a DSLR reference. Performance was assessed using CIEDE2000 color differences ($\Delta E$) and Intraclass Correlation Coefficients (ICC) for Melanin Index, Erythema Index, and Individual Typology Angle (ITA).
@@ 1 Introduction
The standard solution—linear Color Correction Matrix (CCM) normalization—transforms device RGB values toward a reference standard by minimizing the Euclidean distance in color space.[Hong2001, Cheung2004] This approach optimizes for perceptual color fidelity, as measured by the CIEDE2000 color difference formula ($\Delta E_{00}$).[Sharma2005, Luo2001] The clinical acceptability threshold of $\Delta E<2$ corresponds to the just-noticeable difference in human color perception.[Melgosa1997]
@@
Prior research has validated smartphone telemedicine for specific conditions,[Lamel2012, VanoGalvan2011] developed multi-device domain adaptation for other imaging modalities,[Chen2019, Tellez2019, Guan2021] and established RGB-based chromophore quantification methods.[Tsumura2003, Stamatas2004] However, these studies predominantly utilized datasets composed of Fitzpatrick I–II skin types,[Daneshjou2022, Groh2021] and none systematically tested whether colorimetric calibration guarantees clinical biomarker preservation.
This study addresses this critical gap. Using 43,424 images from 965 Korean subjects (Fitzpatrick III–IV), we test the hypothesis that standard colorimetric calibration ensures clinical reliability. Our investigation reveals three key findings:
@@
    Color-Clinical Decoupling: We introduce and systematically investigate “color-clinical decoupling”—the phenomenon where perceptually accurate calibration ($\Delta E<2.3$) fails to preserve clinical biomarkers. While prior work has identified ITA measurement limitations including sensitivity to illuminants,[Jung2024] this is the first systematic demonstration that standard colorimetric calibration creates false clinical confidence, with ITA showing poor agreement (ICC =0.46=0.46) despite excellent color fidelity. We provide mathematical derivation explaining this phenomenon through the differential sensitivity of ratio-based formulas to channel-specific errors.
@@ 2.1 Optical Properties of Human Skin
Human skin functions as a multilayered turbid medium whose appearance is governed by the interplay of absorption and scattering within the epidermis and dermis.[Tuchin2000, Jacques2013] The two principal chromophores are melanin, concentrated in the basal layer of the epidermis, and hemoglobin, distributed within the dermal vasculature.[Anderson1981] Melanin exhibits broadband absorption across the visible spectrum, decreasing monotonically from 300 to 700 nm. Oxyhemoglobin displays characteristic absorption peaks at 542 nm and 577 nm, and blood stasis patterns further modulate perceived skin pigmentation.[Zonios2001, Stamatas2004b]
The spectral reflectance of skin results from a complex light transport process involving specular reflection at the stratum corneum, volumetric scattering within the dermis, and wavelength-dependent absorption at varying depths.[Meglinski2002, Baranoski2004] RGB sensors integrate this continuous spectrum into three broad channels, irreversibly discarding the fine spectral detail required to deconvolve individual chromophore contributions—a limitation arising from metamerism.
@@ 2.2 Device-Dependent Color Rendering
Modern digital cameras exhibit substantial variability in color reproduction due to differences in sensor design and ISP pipelines.[Karaimer2016] Three primary sources of device dependence exist: (1) spectral sensitivity functions of color filter arrays vary across manufacturers;[Hong2001, Chakrabarti2014] (2) ISP pipelines apply nonlinear transformations including gamma correction, tone mapping, and “memory color” optimization;[Karaimer2016, Cheung2004] and (3) the sRGB color space provides limited gamut and introduces quantization errors in dark tones.
@@ 2.3 Clinical Skin Color Indices
Quantitative dermatology employs several derived indices to characterize skin pigmentation. The Individual Typology Angle (ITA) provides an objective correlate of Fitzpatrick skin phototype.[Chardon1991, Wilkes2015, DelBino2006] The Melanin Index, computed from $L^{*}$ using logarithmic transformations, correlates with epidermal melanin concentration.[Chardon1991, Stamatas2004] The Erythema Index, derived from the $a^{*}$ channel, reflects dermal hemoglobin content.[Fullerton1996, Takiwaki1998] Critically, while device validation studies have assessed inter-device agreement for individual indices, no prior work has systematically compared how different clinical metrics—with their distinct mathematical structures—respond to identical calibration procedures. This gap is significant: a calibration optimized for one index may fail another.
@@ 2.4 The Implicit Assumption of Colorimetric Calibration
For typical skin colors in our dataset ($L^{*}\approx 75$, $b^{*}\approx 15$), equation (5) yields $|\partial\text{ITA}/\partial b^{*}|\approx 0.03$ rad/unit $\approx 1.7$°/unit. A seemingly small 1-unit error in $b^{*}$ produces a 1.7° shift in ITA—sufficient to cross Fitzpatrick skin type boundaries—while contributing only $\sim$1 to the $\Delta E$ calculation. Notably, perceptibility thresholds differ across CIELAB axes, with the $b^{*}$ (blue-yellow) axis generally exhibiting higher tolerance thresholds than $L^{*}$ or $a^{*}$, meaning $b^{*}$ errors are both less perceptible and more clinically consequential.
The $b^{*}$ channel is derived primarily from the blue sensor. Due to the shallow absorption depth of blue photons in silicon and the spectral deficiency of common indoor illuminants, the effective signal-to-noise ratio of the blue channel is significantly lower than that of the green channel.[Nakamura2005, Fossum1997] Shot noise scales as $\sigma\propto 1/\sqrt{QE}$, making blue the noisiest channel. This noise propagates through sRGB-to-CIELAB transformation and is amplified by the ITA formula’s sensitivity to $b^{*}$.
@@ 2.5 Dataset Bias in Dermatological AI
While underrepresentation of darker skin types in dermatological AI is well-documented,[Kinyanjui2020, Groh2021, Adamson2018] our dataset of 43,424 images from 965 Korean subjects (Fitzpatrick III–IV) represents the largest cohort specifically designed to evaluate color-clinical decoupling in this underrepresented population, providing systematic characterization of device-dependent color rendering.
@@ 3.1.2 Imaging Devices
    DSLR (Reference): Professional-grade camera under controlled dark room lighting, capturing 7 viewing angles (front, front-top, front-bottom, left 15°, left 30°, right 15°, right 30°)
@@
    Tablet: Consumer-grade tablet camera, capturing 3 viewing angles (front, left, right)
@@
    Smartphone: Consumer-grade smartphone camera, capturing 3 viewing angles (front, left, right)
The analyzed dataset comprises 43,424 samples: DSLR (33,774 samples across 7 angles), tablet (4,825 front-facing samples), and smartphone (4,825 front-facing samples). One DSLR sample was excluded due to invalid bounding box coordinates that prevented reliable color extraction. Although the source dataset includes 3 angles for consumer devices (totaling 14,475 samples per device), only the front-facing imagery was included in this study to maintain focus on cross-device consistency. For cross-device agreement analysis, these 14,475 matched front-facing samples (4,825 per device) ensured geometric alignment across all three platforms.
@@ 4.1 Raw Device Performance: Clinical Unusability of Uncalibrated Images
Table 1 summarizes device characteristics. Consumer devices exhibited systematic underexposure relative to DSLR ($L^{*}=72.7$$77.6$ vs. $81.3$) and compressed $b^{*}$ values (8.968.9610.1210.12 vs. $17.59$), as visualized in Figure 1 and Figure 2.
@@
Figure 1: RGB and CIELAB color distributions across three imaging devices (DSLR, tablet, smartphone) for Korean skin samples (total n=43,424; cross-device comparison subset n=14,475). (A) RGB channel distributions showing device-specific color rendering characteristics. (B) CIELAB $L^{*}a^{*}b^{*}$ distributions demonstrating lightness and chromaticity variations across devices.
Figure 2: Signal-to-noise ratio (SNR) comparison across devices calculated from repeated measurements at multiple capture angles. Higher SNR values indicate better measurement precision and lower random noise. Error bars represent standard deviation across facial regions (forehead, left cheek, right cheek, chin, glabella).
Figure 3: Bland-Altman plots comparing $L^{*}$, $a^{*}$, and $b^{*}$ measurements between consumer devices (tablet, smartphone) and reference DSLR. Solid horizontal lines indicate mean bias; dashed lines show 95% limits of agreement ($\pm$1.96 SD).
Figure 4: CIEDE2000 color difference ($\Delta E$) distributions by facial region and source device. The vertical dashed line indicates the clinically perceptible threshold ($\Delta E=2.0$). The chin region exhibits the largest color differences, while cheek regions show better inter-device agreement.
@@ 4.2 CCM Normalization: Colorimetric Success
Linear CCM substantially reduced color differences (Table 3 and Figure 5): 77.4% improvement for tablet ($\Delta E$: 8.33 $\rightarrow$ 1.88) and 67.4% for smartphone (7.02 $\rightarrow$ 2.29). Post-normalization median values approached clinical acceptability (1.66 and 2.09 respectively).
@@
Figure 5: Color correction matrix (CCM) normalization results showing $\Delta E$ distributions before and after calibration. Post-normalization $\Delta E$ distributions demonstrate significant improvement, with linear CCM trained using 5-fold cross-validation.
@@ 4.3 The Decoupling: Clinical Index Failure Despite Colorimetric Success
Despite effective colorimetric normalization, clinical index agreement varied dramatically (Table 4, Figure 6, and Figure 7). While Melanin Index achieved good agreement (ICC =0.767=0.767), ITA showed poor agreement (ICC =0.455=0.455). This represents the central finding: perceptual color accuracy does not guarantee clinical biomarker reliability.
@@
Figure 6: Comparison of clinical skin color indices across devices after CCM normalization. (A) Melanin Index distributions. (B) Erythema Index distributions. (C) Individual Typology Angle (ITA) for Fitzpatrick skin type classification.
Figure 7: Pearson correlation plots comparing clinical indices between consumer devices and DSLR reference. Each point represents a matched sample (same subject, same facial region). Diagonal dashed line indicates perfect agreement ($r=1.0$).
@@ 4.4 Sources of Variance: Implications for Calibration Protocol Design
ANOVA revealed a striking hierarchy of variance sources (Table 5 and Figure 8). Facial anatomical region accounted for 25.2% of variance ($\eta^{2}=0.252$)—3.6 times greater than device hardware effects ($\eta^{2}=0.070$).
@@
Figure 8: Stratified analysis of pre-normalization color differences ($\Delta E$) across demographic factors. (A) Age groups. (B) Gender. (C) Fitzpatrick skin type. (D) Facial region. One-way ANOVA revealed significant effects for all factors, with facial region showing the largest effect size.
@@ 5.1 The Color-Clinical Decoupling Phenomenon
The central finding of this study is the non-linear decoupling between colorimetric accuracy and clinical biomarker reliability. Despite achieving near-clinical color fidelity ($\Delta E<2.3$), the ITA index exhibited poor inter-device agreement (ICC =0.46=0.46). This finding challenges the fundamental assumption underlying smartphone dermatology: that “looking correct” guarantees “diagnosing correctly.”
@@
Critically, this sensitivity is phenotype-dependent. The ITA sensitivity formula $|\partial\text{ITA}/\partial b^{*}|=|L^{*}-50|/[(L^{*}-50)^{2}+(b^{*})^{2}]$ reveals that darker skin types—characterized by lower $L^{*}$ values closer to 50 and lower $b^{*}$ values—experience dramatically amplified error sensitivity. For Type V skin ($L^{*}\approx 60$, $b^{*}\approx 10$), the sensitivity increases to $\sim$2.9°/unit; for Type VI skin ($L^{*}\approx 40$, $b^{*}\approx 5$), it reaches $\sim$4.6°/unit. A 2-unit calibration residual in $b^{*}$—perceptually negligible—produces a 3.4° shift in ITA for light skin but a 9.2° shift for dark skin. This represents a “metrological bias”: current calibration protocols, by optimizing for average $\Delta E$, do not suppress $b^{*}$ errors with the aggression required for accurate assessment of darker skin tones. Table 6 summarizes this phenotype-dependent vulnerability.
The blue sensor’s inherently low quantum efficiency produces elevated shot noise that propagates directly to $b^{*}$. Linear CCM, optimized for overall color fidelity, cannot selectively suppress this critical error source. The result is “color-clinical decoupling”: colorimetric success masking clinical failure.
@@
Table 6: Phenotype-dependent ITA sensitivity to $b^{*}$ channel errors. Sensitivity calculated from $|\partial\text{ITA}/\partial b^{*}|=|L^{*}-50|/[(L^{*}-50)^{2}+(b^{*})^{2}]$. The “2-unit Error Impact” shows the ITA shift from a typical calibration residual, demonstrating disproportionate effects on darker skin types.
Skin Type $L^{*}$ $b^{*}$ Sensitivity (°/unit) 2-unit Error Impact
Type II/III (Light) 75 15 1.68 3.4°
Type IV (Intermediate) 65 15 1.92 3.8°
Type V (Brown) 60 10 2.86 5.7°
Type VI (Dark) 40 5 4.58 9.2°
@@ 5.2 Why Single-Patch Calibration Fails
This “anatomical dominance” arises from multiple physiological sources. The T-zone (forehead and glabella) exhibits higher sebaceous gland density than cheek regions, resulting in elevated sebum secretion that alters surface optical properties.[Man2009] Sebum functions as an index-matching fluid ($n\approx 1.46$), reducing the refractive index mismatch at the stratum corneum ($n\approx 1.55$) via the optical clearing effect.[Tuchin2000, Ding2006] This reduction in refractive index discontinuity decreases surface scattering, allowing more photons to penetrate deeper into the epidermis and dermis where they interact with melanin and hemoglobin—producing “richer” or “darker” apparent skin color (lower $L^{*}$, higher chroma). Simultaneously, the smooth liquid sebum surface produces intense specular highlights under directional illumination.[Baranoski2004] These regional optical variations produce angle-dependent color shifts that a single CCM cannot correct.
Facial skin thickness varies substantially: the periorbital region is among the thinnest (0.5–1.0 mm), while chin and forehead are thicker (1.5–2.0 mm).[Chopra2015] Since skin color arises from chromophore absorption at specific depths, thickness variations alter the optical path length and relative contributions of melanin versus hemoglobin.